OpenStudy (anonymous):

PLEASE HELP WITH WORD PROBLEM??? A cup of coffee was made at a temperature of 90C and cools according to Newton's law of cooling. The room temperature is 30C. If the temperature of the coffee 20 minutes after being made was 40 C When was the temperature of the coffee 80C Time = how many minutes??

OpenStudy (campbell_st):

Is the formula something like \[T_{n} = T_{o} + e^{kt}\]

OpenStudy (anonymous):

i believe so i know after 10 minutes the temp. is 54 degrees

OpenStudy (campbell_st):

something is wrong in the formula... have you been given a formula for this..?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i haven't been given a formula for this..

OpenStudy (campbell_st):

ok... I've just done some reading online it looks like \[T = e^{kt + C} + 30\] so start by finding C, which uses the initial values t= 0 and T = 90

OpenStudy (anonymous):

hold on let me check

OpenStudy (anonymous):

for some reason its not correct

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

oops... found another mistake... if you give me a minute I'll correct everything...

OpenStudy (anonymous):

ok thanks!

OpenStudy (campbell_st):

ok so the model is \[T = e^{kt + C} + 30\] start by using the initial conditions t = 0 and T = 90 to find C \[C = \ln(90 -30)\] so C = 4.09434 the model now becomes \[T = e^{kt + 4.09434} + 30\] next use t = 20 and T = 40 to find the value of k \[40 = e^{20k + 4.09434} + 30\] which simplifies to \[k = \frac{\ln(10) - 4.09434}{20}\] so k = -0.089588 so the final model is \[T = e^{-0.089588t + 4.09434} + 30\] now you need to find t, when T = 80 \[80 = e^{-0.089588 + 4.09434} + 30\] solving for t you get \[t = \frac{\ln(50) - 4.09434}{-0.089588}\] so t = 2.03 mins or the coffee cools to 80 in 2.03 mins

OpenStudy (campbell_st):

I can't tell from the question is your answer should be 2 mins or 2.03 mins..

OpenStudy (anonymous):

2.03 is correct, your sooo amazing!!! :) Can you help me with another question??

OpenStudy (campbell_st):

lol... sorry I have to go to school....